# On toric foliated pairs

### v1

@@ 6. Appendix: Toric projective bundles
Since $X$ is a $\mathbb{P}^{r}$-bundle over $Y$, we can write $X=\mathbb{P}_{Y}(\mathcal{E})$ for some vector bundle $\mathcal{E}$ on $Y$. We take a torus invariant Cartier divisor $H$ on $X$ such that $\mathcal{O}_{X}(H)\simeq\mathcal{O}_{\mathbb{P}_{Y}(\mathcal{E})}(1)$. Then we have $\varphi_{*}\mathcal{O}_{X}(H)\simeq\mathcal{E}$. Thus, by replacing $\mathcal{E}$ with $\varphi_{*}\mathcal{O}_{X}(H)$, we may assume that $\mathcal{E}$ is a toric vector bundle on $Y$, that is, the torus action on $Y$ lifts to an action on $\mathcal{E}$ and it is linear on the fibers. Let $U$ be any affine toric open subset $U$ of $Y$. Then it is not difficult to see that $\mathcal{E}|_{U}$ is isomorphic to $\mathcal{O}^{\oplus r+1}_{U}$ as a toric vector bundle on $U$ (see, for example, [Py, Proposition 2.2]). Therefore, the restriction of $\varphi\colon X\to Y$ to $U$ is isomorphic to the second projection $\mathbb{P}^{r}\times U\to U$. Let $h\colon(N,\Sigma)\to(N^{\prime},\Sigma^{\prime})$ be a map of fans corresponding to $\varphi\colon X\to Y$. Let $N^{\prime\prime}$ be the kernel of $h\colon N\to N^{\prime}$. Without loss of generality, we may assume that $N=N^{\prime}\oplus N^{\prime\prime}$. We fix a $\mathbb{Z}$-basis $\{n^{\prime\prime}_{1},\ldots,n^{\prime\prime}_{r}\}$ of $N^{\prime\prime}$. Since $\varphi\colon X\to Y$ is isomorphic to the second projection $\mathbb{P}^{r}\times U\to U$ for any affine toric open subset $U$ of $Y$, we can lift any cone $\sigma^{\prime}\in\Sigma^{\prime}$ to a cone $\sigma\in\Sigma$. Hence we can find $\Sigma^{\prime}$-linear support functions $h_{1},\ldots,h_{r}$ such that the map $N^{\prime}_{\mathbb{R}}\to N_{\mathbb{R}}=N^{\prime}_{\mathbb{R}}\oplus N^{\prime\prime}_{\mathbb{R}}$ given by $y\mapsto\left(y,-\sum_{i=1}^{r}h_{i}(y)n^{\prime\prime}_{i}\right)$ defines the desired lifts of cones. Let $\mathcal{L}_{i}$ be the line bundle on $Y$ defined by the $\Sigma^{\prime}$-linear support function $h_{i}$ for every $i$. Then, by construction, we can check that $X\simeq\mathbb{P}_{Y}(\mathcal{O}_{Y}\oplus\mathcal{L}_{1}\oplus\cdots\oplus\mathcal{L}_{r})$ and $\varphi\colon X\to Y$ is isomorphic to the projection $\pi\colon\mathbb{P}_{Y}(\mathcal{O}_{Y}\oplus\mathcal{L}_{1}\oplus\cdots\oplus\mathcal{L}_{r})\to Y$. We finish the proof. ∎

### v3

@@ 6. Appendix: Toric projective bundles
Since $X$ is a $\mathbb{P}^{r}$-bundle over $Y$, we can write $X=\mathbb{P}_{Y}(\mathcal{E})$ for some vector bundle $\mathcal{E}$ on $Y$. We take a torus invariant Cartier divisor $H$ on $X$ such that $\mathcal{O}_{X}(H)\simeq\mathcal{O}_{\mathbb{P}_{Y}(\mathcal{E})}(1)$. Then we have $\varphi_{*}\mathcal{O}_{X}(H)\simeq\mathcal{E}$. Thus, by replacing $\mathcal{E}$ with $\varphi_{*}\mathcal{O}_{X}(H)$, we may assume that $\mathcal{E}$ is a toric vector bundle on $Y$, that is, the torus action on $Y$ lifts to an action on $\mathcal{E}$ and it is linear on the fibers. Let $U$ be any affine toric open subset $U$ of $Y$. Then it is not difficult to see that $\mathcal{E}|_{U}$ is isomorphic to $\mathcal{O}^{\oplus r+1}_{U}$ as a toric vector bundle on $U$ (see, for example, [Pay, Proposition 2.2]). Therefore, the restriction of $\varphi\colon X\to Y$ to $U$ is isomorphic to the second projection $\mathbb{P}^{r}\times U\to U$. Let $h\colon(N,\Sigma)\to(N^{\prime},\Sigma^{\prime})$ be a map of fans corresponding to $\varphi\colon X\to Y$. Let $N^{\prime\prime}$ be the kernel of $h\colon N\to N^{\prime}$. Without loss of generality, we may assume that $N=N^{\prime\prime}\oplus N^{\prime}$. We fix a $\mathbb{Z}$-basis $\{n^{\prime\prime}_{1},\ldots,n^{\prime\prime}_{r}\}$ of $N^{\prime\prime}$. Since $\varphi\colon X\to Y$ is isomorphic to the second projection $\mathbb{P}^{r}\times U\to U$ for any affine toric open subset $U$ of $Y$, we can lift any cone $\sigma^{\prime}\in\Sigma^{\prime}$ to a cone $\sigma\in\Sigma$. Hence we can find $\Sigma^{\prime}$-linear support functions $h_{1},\ldots,h_{r}$ such that the map $N^{\prime}_{\mathbb{R}}\to N_{\mathbb{R}}=N^{\prime\prime}_{\mathbb{R}}\oplus N^{\prime}_{\mathbb{R}}$ given by $y\mapsto\left(\sum_{i=1}^{r}h_{i}(y)n^{\prime\prime}_{i},y\right)$ defines the desired lifts of cones. Let $\mathcal{L}_{i}$ be the line bundle on $Y$ defined by the $\Sigma^{\prime}$-linear support function $h_{i}$ for every $i$. Then, by construction, we can check that $X\simeq\mathbb{P}_{Y}(\mathcal{O}_{Y}\oplus\mathcal{L}_{1}\oplus\cdots\oplus\mathcal{L}_{r})$ and $\varphi\colon X\to Y$ is isomorphic to the projection $\pi\colon\mathbb{P}_{Y}(\mathcal{O}_{Y}\oplus\mathcal{L}_{1}\oplus\cdots\oplus\mathcal{L}_{r})\to Y$ (see Remark 6.2 below). We finish the proof. ∎
Remark 6.2 (see [Par, p.124, Remark.(2)]).
The minus sign in [O2, p.59] needs to be deleted.